# Understanding The Concerns and Choices of The Public When Using Large Language Models for Healthcare

### v1

@@ 5.4. Limitation
There are several limitations of our research. Firstly, almost all our interview participants were between the ages of 20 and 30. Our interview sample was also limited to Asian and Black people. As a result, we were unable to learn about the experiences and challenges of other age groups and races when using LLMs for healthcare. For example, Middle-aged and elderly people may have higher and more complex needs for healthcare. Future research should consider more diverse demographic groups.
@@ 6. Conclusion
We use a mix of surveys and interviews to study the concerns and choices of the public when using large language models for healthcare. Compared with prior research on public healthcare and large language models, our research has the following three contributions: first, through the surveys, we learned about the different ways the public uses to obtain healthcare information from LLMs; second, through both surveys and interviews, we learned about different motivations of the public for using LLMs for healthcare; third, we learned about the public’s value judgments when facing some of the controversial issues, such as misinformation and doctor-patient relationships. We concluded by providing recommendations for the public to better use LLMs to search for health information and for doctors to more ethically use LLMs for diagnosis. We also pointed out future research directions in LLMs for public health.

### v2

@@ 5.4. Limitation
There are several limitations of our research. Firstly, limited by information channels and inclusion criteria, almost all our interview participants were between the ages of 20 and 30, and lots of them were students. Most of the survey participants are between the ages of 20 and 39, although this is roughly consistent with the age group of Chatgpt users (None, 2017c). Our interview sample was also limited to Asian and Black people. As a result, we were unable to learn about the experiences and challenges of other age groups and races when using LLMs for healthcare. For example, Middle-aged and elderly people may have higher and more complex needs for healthcare. We hope future research should consider more diverse demographic groups, especially the middle-aged and elderly.
@@ 6. Conclusion
We use a mix of surveys and interviews to study the concerns and choices of the public when using large language models for healthcare. Compared with prior research on public healthcare and large language models, our research has the following three contributions: first, through the surveys, we learned about the different ways the public uses to obtain healthcare information from LLMs; second, through both surveys and interviews, we learned about different motivations of the public for using LLMs for healthcare; third, we learned about the public’s value judgments when facing some of the controversial issues, such as misinformation and doctor-patient relationships. We concluded by providing recommendations for the public to use LLMs better to search for health information and for doctors to use LLMs for diagnosis more ethically. We also pointed out future research directions in LLMs for public health.